Brandtly Industries invests a large sum of money in R&D; as a result, it retains and reinvests all of its earnings. In other words, Brandtly does not pay any dividends, and it has no plans to pay dividends in the near future. A major pension fund is interested in purchasing Brandtly's stock. The pension fund manager has estimated Brandtly's free cash flows for the next 4 years as follows: $4 million, $7 million, $8 million, and $16 million. After the fourth year, free cash flow is projected to grow at a constant 6%. Brandtly's WACC is 14%, the market value of its debt and preferred stock totals $40 million, the firm has $14 million in non-operating assets, and it has 20 million shares of common stock outstanding.

  1. What is the present value of the free cash flows projected during the next 4 years? Do not round intermediate calculations. Round your answer to the nearest dollar. Write out your answers completely. For example, 13 million should be entered as 13,000,000.
    $   

  2. What is the firm's horizon, or continuing, value? Round your answer to the nearest dollar. Write out your answers completely. For example, 13 million should be entered as 13,000,000.
    $   

  3. What is the market value of the company's operations? Do not round intermediate calculations. Round your answer to the nearest dollar. Write out your answers completely. For example, 13 million should be entered as 13,000,000.
    $   

    What is the firm's total market value today? Do not round intermediate calculations. Round your answer to the nearest dollar. Write out your answers completely. For example, 13 million should be entered as 13,000,000.
    $   

  4. What is an estimate of Brandtly's price per share? Do not round intermediate calculations. Round your answer to the nearest cent.
    $   
Answer #1

a.Present value = Future value/(1+Interest rate)^Number of years

Present value of free cash flows of next 4 years = 4,000,000/(1.14) + 7,000,000/(1.14)^2 + 8,000,000/(1.14)^3 + 16,000,000/(1.14)^4

= $23,768,101.1965

i.e. $23,768,101

b.Horizon value = Cash flow in year 5/(Required return – growth rate)

= 16,000,000(1+6%)/(14%-6%)

= $212,000,000

c.Market value of operations = Present value of all future free cash flows

= 23,768,101 + 212,000,000/(1.14)^4

= $149,289,119.80

i.e. $149,289,120

Total market value = Market value of operations + non operating assets

= 149,289,120 + 14,000,000

= $163,289,120

g.Price per share = (Value of firm – value of debt and preferred stock)/Number of shares

= (163,289,120-40,000,000)/20,000,000

= $6.1645 per share

i.e. $6.16 per share